OpenStudy (anonymous):

I Give Metals. Which of the following is an equivalent representation of 3^-3 ?

OpenStudy (anonymous):

 @texaschic101

OpenStudy (anonymous):

a 1/27 b. 1/9 c. 1/64 d. 1/16

OpenStudy (texaschic101):

a^-n = 1/a^n 3^-3 = 1/(3^3) or 1/27

OpenStudy (anonymous):

can you help me with more

OpenStudy (texaschic101):

I can try..but I am not the best with exponents

OpenStudy (anonymous):

Which of the following best describes the solution to the equation below? 14x + 15 = 7x + 32

OpenStudy (texaschic101):

little hint....when moving something across the equal sign, you just have to change the sign. 14x + 15 = 7x + 32 --- gonna move the 15 and the 7 across the equal sign, so I have to change their sign... 14x - 7x = 32 - 15 -- now combine terms 7x = 17 -- now divide both sides by 7 x = 17/7

OpenStudy (anonymous):

A exactly one real solution,x= 2 3/7 B. no real solutions C. exactly one real solution, x=17 D. infinite real solutions

OpenStudy (texaschic101):

exactly one real solution....17/7 = 2 3/7....so A

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

x/5 - 8 = 7

OpenStudy (anonymous):

A x=90 B. x= -5 C.x=60 D. 75

OpenStudy (anonymous):

D x=75 i mean

OpenStudy (texaschic101):

x/5 - 8 = 7....add 8 to both sides x/5 = 15....now multiply both sides by 5 x = 75

OpenStudy (anonymous):

Find an expression equivalent to the one shown below. 12^7*12^negative 2

OpenStudy (anonymous):